OpenStudy (anonymous):

2(x+6)-10=12+4(x-1) explain the process for solving equation Please!

OpenStudy (anonymous):

first, you need to use the distributive property: \[a(b+c) =ab + ac\] to simplify: \[2x+12 - 10 = 12 + 4x-4\]

OpenStudy (anonymous):

then you should combine the like terms to simplify again: \[2x + 2 = 8 + 4x\]

OpenStudy (anonymous):

now that the equation is simplified, you should work on isolating the x

OpenStudy (anonymous):

subtract 2x from both sides to get: \[2 = 8+2x\] subtract 8 from both sides to get: \[-6 = 2x\]

OpenStudy (anonymous):

now that you have an x and a coefficient, divide by the coefficient: \[-3 = x\]

OpenStudy (anonymous):

you can check by plugging x back into the equation: \[2(-3 + 6) - 10 = 2(3) - 10 = 6-10 = -4\] \[12+4(-3 - 1) = 12 + 4(-4) = 12 - 16 = -4\] both sides are equal, so this is the value of x